Imagine that Canada produces only three goods: apples, bananas, and carrots. The quantities produced and the prices of the three goods are listed below:

Prices (S) 2.00 1.50 1.00 Goods Quantities produced es 12 20 Bananas Instructions: Round your answers to the nearest dollar, a. Canadian GDP is$ b. Suppose a drought hits the province of British Columbia. This drought causes the quantity of apples produced to fall to 6 56 50 Assuming that all prices remain constant, the new Canadian GDP is S Instructions: Round your answer to the nearest cent c. Assume, once again, that the quantities produced and the prices of the three goods are as listed in the table. Now, given this situation, carrot sellers decide that the price of carrots is too low, so they ogree to raise the price. the Canadian GDP s $80.00.the new price of carrots ss 22 per carrot 2.2 per carrot. References eBook & Resources
Answer #1

GDP is the market value of all final goods and services produced in a country in a year. Canada only produces three goods. The GDP can be calculated by adding the value of all three goods produced there.

GDP = Price of apples*Quantity of apples + Price of bananas*Quantity of bananas + Price of carrots*Quantity of carrots

a) Canada's GDP = (9*2) + (12*1.5) + (20*1)

= 18 + 18 + 20 = $56

b) Due to drought , the quantity of apples produced declines to 6.

GDP = (6*2) + (12*1.5) + (20*1) = $50

c) Carrot sellers have raised the price and the resulting GDP is $80. Suppose the new price is x.

80 = (9*2) + (12*1.5) + (20*x)

80 - 18 - 18 = 20x

44 = 20x

x = $2.2

That means that price of carrots have been raised to $2.2 per carrot.